OpenStudy (anonymous):

A ball is thrown in the air. the relationship between the time the ball is in the air in seconds(t) and the height of a ball in feet above the ground(h) is represented by h=-16t^2+46t+6.How many seconds will it take for the ball to hit the ground?

OpenStudy (anonymous):

when it will hit ground h=0 solve -16t^2+46t+6=0

OpenStudy (anonymous):

I got multiple choices A) 1/8 B)3/8 C)3 D)6 just need the steps to understand

OpenStudy (anonymous):

When it will hit the ground translates to the fact that your height will be 0 ft so that means h=0 so you want to solve for t in the equation h=-16t^2+46t+6 when h=0: 0=-16t^2+46t+6 do you know where to begin?

OpenStudy (anonymous):

maybe we have to factor it?!

OpenStudy (anonymous):

yes for a quadratic you can factor or use the quadratic equation

OpenStudy (anonymous):

ok you need to solve -16t^2+46t+6=0 take -2 common -2(8t^2-23t-3)=0 break mid term to have factors . -2(8t^2 -24t+t-3)=0 -2(8t(t-3)+1(t-3))=0 (8t+1)(t-3)=0 t=3 t=-1/8

OpenStudy (anonymous):

ok thanks that really helped